### 15 Dec 2023 v2

# Chat-3D v2: Bridging 3D Scene and Large Language Models with Object Identifiers

Abstract
Recent research has evidenced the significant potentials of Large Language Models (LLMs) in handling challenging tasks within 3D scenes. However, current models are constrained to addressing object-centric tasks, where each question-answer pair focuses solely on an individual object. In real-world applications, users may pose queries involving multiple objects or expect for answers that precisely reference various objects. We introduce the use of object identifiers to freely reference objects during a conversation. While this solution appears straightforward, it presents two main challenges: 1) How to establish a reliable one-to-one correspondence between each object and its identifier? 2) How to incorporate complex spatial relationships among dozens of objects into the embedding space of the LLM? To address these challenges, we propose a two-stage alignment method, which involves learning an attribute-aware token and a relation-aware token for each object. These tokens capture the object’s attributes and spatial relationships with surrounding objects in the 3D scene. Once the alignment is established, we can fine-tune our model on various downstream tasks using instruction tuning. Experiments conducted on traditional datasets like ScanQA, ScanRefer, and Nr3D/Sr3D showcase the effectiveness of our proposed method. Additionally, we create a 3D scene captioning dataset annotated with rich object identifiers, with the assistant of GPT-4. This dataset aims to further explore the capability of object identifiers in effective object referencing and precise scene understanding. [^1] Code will be released at https://github.com/Chat-3D/Chat-3D-v2.
Figure 1: An example of using object identifiers in the conversation.
1 Introduction
Researchers have recently shown a strong interest in general-purpose Multi-modal Large Language Models (MLLMs) [21, 48, 11, 13, 12]. These models, through establishing alignment between language and various modalities such as images, audio, and point clouds, gain the ability to comprehend multi-modal content and tackle diverse tasks. Although methods like Imagebind-LLM [13] and Point-LLM [12] have been developed to align 3D point clouds with LLMs, adapting them to real-world 3D scenes remains challenging. This is primarily due to the intricate spatial relationships among numerous objects and the limited availability of 3D scene-text data used for effective alignment.
3D-LLM [14] and Chat-3D [33] take the initial efforts to incorporate 3D scenes into LLMs. They effectively address a wide range of 3D-language scene understanding tasks, highlighting the potential of LLMs in comprehending the complexities of real 3D scenes. 3D-LLM primarily relies on translating 3D features into the space of 2D Vision-Language Models (VLMs) [19, 2]. However, handling the intricate spatial relationships inherent in 3D scenes proves challenging for the 2D vision encoder of the VLM. On the other hand, Chat-3D offers a solution by directly mapping 3D features into the LLM space and employing a relation module to capture spatial relationships, which demonstrates impressive conversational capabilities within 3D scenes. Nevertheless, it requires specifying a target object for the conversation, which limits the range of dialogue content and poses challenges in addressing complex tasks in real-world applications.
When referring to a specific object within a real 3D scene, such as the chair located inside the red box in Figure 1, it is often necessary to provide a detailed description to distinguish it from other similar chairs nearby. For example, you can specify its position as follows: “the chair located at the southwest corner of the rightmost table.” However, using such lengthy descriptions not only complicates the interaction between the user and the assistant but can also introduce ambiguity. This ambiguity may confuse the language model, especially when terms like “rightmost” are open to interpretation depending on the observer’s perspective. This challenge becomes even more pronounced when dealing with complex tasks involving multiple objects that are not easily described using natural language.
We propose assigning unique identifiers for every object instances within the 3D scene. These instances can be easily extracted by advanced and reliable 3D instance segmentation models [42, 16, 29], benefiting from the mature development of the 3D segmentation field. As depicted in Figure 1, the chair, and two trash cans are labeled as “obj13”, “obj23”, and “obj32”, respectively. These identifiers can be utilized in conversations, facilitating clearer communication between the user and the assistant. Firstly, it allows the user to unambiguously reference specific objects using these predefined identifiers, eliminating confusion arising from vague descriptions. Secondly, the language assistant can respond using these specific identifiers, which can also be visualized within the original scene, aiding in addressing more complex tasks and bringing a superior user experience.
To fully harness the advantages of object identifiers, we must confront two substantial challenges: 1) In contrast to previous methods that involve a fixed number of visual tokens in the language model, our approach necessitates the establishment of a one-to-one correspondence between each object token and its associated identifier. This becomes particularly crucial when dealing with object tokens originating from 3D space, rather than conventional words like “chair.” 2) It is a challenging task to tackle the complex spatial relationships among dozens of objects within a 3D scene. While we can rely on the inherent reasoning capabilities of the Language Model, there exists a largely unexplored potential to incorporate more 3D spatial information into the language model, thus enhancing its spatial understanding.
To address the challenges mentioned above, we propose to learn both attribute-aware and relation-aware tokens for each object, which are then combined with unique identifiers. We devise a step-by-step aligning strategy, which consists of object-level alignment and scene-level alignment, aimed at mapping object features from the 3D encoder space into the language model’s embedding space. In the object-level alignment phase, we initially learn a 3D-language projection to integrate the 3D object feature into an attribute-aware token for each object. Moving on to the scene-level alignment phase, we design a relation module to incorporate spatial information into a relation-aware token for each object, in conjunction with the identifier and the attribute-aware token. Once alignment is established, we can fine-tune the model on various downstream tasks, enabling it to adapt to the target answer formats and ultimately achieve improved performance.
Another challenge lies in the limitations of the existing 3D-language scene understanding datasets [4, 1, 3], in which one language query only focuses on a single or few objects. Thus, models tested on these datasets are not able to demonstrate a comprehensive understanding of the entire 3D scene. To tackle this, we create an identifier-rich scene captioning dataset with the assist of GPT-4 to further explore our model’s scene understanding and reasoning abilities.
We also evaluate our model on traditional 3D-language downstream datasets, such as ScanQA [3], ScanRefer [4], and Nr3D/Sr3D [1]. Our model outperforms previous 3D large language models and achieve comparable results with traditional supervised baselines.
Our contributions can be summarized as follows:
  • We develop an LLM-based model for 3D scene understanding, equipped with unique object identifiers to facilitate effective object referencing and enhance task-solving capabilities.
  • To establish a one-to-one correspondence between each object and its identifier, we propose integrating 3D objects into the LLM through a two-stage alignment process, which learns attribute-aware and relation-aware tokens for each object to capture the object attributes and complex spatial relationships within the 3D scene.
  • We create an identifier-rich scene captioning dataset, with the assistant of GPT-4, to explore the comprehensive scene understanding and reasoning ability of our model. Experimental results underscore the importance of object identifiers in referencing objects freely and enhancing the overall scene understanding.
  • We also evaluate our model on traditional 3D-language downstream datasets, such as ScanQA, ScanRefer, and Nr3D/Sr3D. Our model outperforms previous 3D large language models and achieve comparable results with supervised baselines.
2 Related Work
2.1 3D-Language Scene Understanding
In the realm of 3D scene understanding, there is a growing focus on leveraging language as background knowledge and query conditions to capture user intentions, given its broad application prospects. This concept is termed “3D-language scene understanding”, covering various downstream tasks: 3D Question Answering [3, 49, 26, 23] necessitates models to answer questions based on 3D scenes; 3D Visual Grounding [4, 15, 5, 31, 45, 32] requires models to localize a target object in the 3D scene according to a language query; 3D Dense Captioning [6, 41, 17] demands proficiency in both localizing and captioning all the objects in the scene. While supervised expert models have been widely explored and demonstrate impressive performance on specific tasks, these models often lack adaptability to other tasks. To devise an unified architecture for all these tasks, an important step is to enable the model’s ability to reference any object in the scene freely, a challenge addressed by our proposed object identifiers. Additionally, we introduce a new 3D scene captioning dataset to better evaluate the model’s scene understanding and reasoning ability.
Figure 2: Overall architecture of our model.
2.2 Multi-modal Large Language Models
Recently, there have been significant advancements in large language models (LLMs) [7, 24, 28, 8], showcasing impressive capabilities in complex reasoning and conversational interactions with humans. Researchers have shown great interest in extending the LLM’s knowledge and abilities to other modalities [20, 21, 46, 48, 13, 12, 14, 33, 38, 18]. Among these, Imagebind-LLM [13] and Point-LLM [12] are the first to integrate 3D modality into LLMs. They achieve this by learning a joint embedding space between 3D point clouds and other modalities, such as image, audio, and text. While effective at the object level, these models face challenges when directly applied to 3D scenes that require understanding complex spatial relationships among objects. 3D-LLM [14] attempts to enhance scene understanding by incorporating positional embedding and learning location tokens. However, they project 3D features into the input space of pre-trained 2D Vision-Language Models (VLMs), encoding the projected 3D feature using a 2D vision encoder before sending it to the language model. This approach poses difficulties for the 2D encoders in capturing the 3D spatial structure and complex relationships among 3D objects. Chat-3D [33] addresses this challenge by directly using 3D scene-text data to align the 3D scene with the LLM, mitigating the problem of limited available data through a pre-alignment stage. However, this model makes a compromise in its architecture design, limiting its ability to focus on only one target object during a conversation. To overcome the shortcomings of previous methods, we propose explicitly combining unique object identifiers with each object’s feature. This approach makes the LLM aware of the existence of each object in the scene and enables it to reference them. Through our proposed alignment strategies, we demonstrate that our model outperforms 3D-LLM using only 3D scene-text data, which is 10$\times$ less than the data used by 3D-LLM, proving the effectiveness of our proposed method.
3D Representation Learning
Recently, numerous efforts have been made to learn discriminative and robust representations for 3D point clouds, which serve as a fundamental visual modality. Approaches such as PointBERT [39], Point-MAE [25], Transformer-OcCo [30], and Point-m2ae [43] employ self-supervised learning techniques to extract meaningful representations of 3D objects from unlabeled point cloud data. Another set of works [36, 22, 44, 10, 35, 34] seeks to extend representation from other modalities to the 3D domain. For example, ULIP [36] and OpenShape [22] construct 3D-image-text triplets to align point clouds within the CLIP [27] representation space. C-MCR [35] and Ex-MCR [34] learn contrastive representations between various modalities, including 3D point clouds. They leverage knowledge from existing MCR spaces to tackle to challenge of lacking paired data. These robust 3D representations effectively capture detailed information about 3D objects. Our approach involves segmenting the 3D scene at the instance level and extracting a set of object features to represent the entire scene.
3 Method
3.1 Overview
We aim to seamlessly integrate individual objects from a 3D scene into a pre-trained Large Language Model (LLM) and enable explicit object referencing using unique identifiers. Our method, depicted in Figure 2, begins by segmenting object instances and assigning each a unique identifier. To establish a one-to-one correspondence between each object instance and its identifier, we pre-train the model in a two-stage alignment process, comprising object-level and scene-level alignment.
In the object-level alignment phase, we first learn a 3D-language projection to integrate the 3D object feature into an attribute-aware token for each object. Moving on to the scene-level alignment phase, we incorporate spatial information using a relation module to create a relation-aware token for each object, along with the identifier and the attribute-aware token. For more detailed information, refer to Section 3.2 and Section 3.3.
After alignment, we can fine-tune the model on various downstream tasks, allowing it to adapt to target answer formats and ultimately achieve improved performance, as discussed in Section 3.4.
3.2 Object-level Alignment
First and foremost, given the thorough development of 3D segmentation techniques [16, 29, 42], it is reasonable to pre-extract all the object instances from the 3D scene using a powerful and reliable instance segmentation model. Each object is assigned a unique identifier in the format ”$\mathrm{obj}[i]$” with $i$ ranging from 0 to $n$, where $n$ represents the total number of objects. Subsequently, we extract object’s 3D features $\mathbf{Z}^{\mathrm{o}}$ using a pre-trained 3D encoder. Additional object attributes $\mathbf{A}\in\mathbb{R}^{n\times 6}$ (RGB color and 3D size of the bounding box) are projected into the object feature space via a linear transformation function $f_{\mathrm{a}}(\cdot)$. This projection results in attribute features denoted as $\mathbf{Z}_{\mathrm{a}}$, which are then concatenated with the object features. Following this, we employ a 3D-language projection layer $f_{\mathrm{o}}(\cdot)$ to map the object features from the 3D encoder feature space to the embedding space of the language model with the dimension of $d$. The final attribute-aware token embedding $\mathbf{F}^{\mathrm{o}}\in\mathbb{R}^{n\times d}$ is obtained by:
$\mathbf{F}^{\mathrm{o}}=f_{\mathrm{o}}([\mathbf{Z}^{\mathrm{o}};f_{\mathrm{a}}(\mathbf{A})]).$ (1)
Attribute-aware Embedding Similarity Initially, we follow the approach of Chat-3D to directly align the attribute-aware token features with the language model in the embedding space. We extract object-level descriptions, such as class name, color, and material, from the ScanRefer dataset. For instance, “a brown wooden shelf.” We compute the mean value of the token embeddings of the description, then maximize the cosine similarity between this mean token embedding and the attribute-aware token embedding. This process establishes an initial connection between the 3D space and the LLM space, creating a smoother learning path for subsequently aligning the object identifiers and incorporating spatial information.
Object-level Question-Answering Next, we associate each attribute-aware token with its corresponding object identifier and input them into the language model. The input for the $i$-th object in the language model is structured as “$\mathrm{obj}[i]:O_{i}$,” where the embedding of the object-level token $O_{i}$ is $\mathbf{F}^{\mathrm{o}}_{i}$ and the embedding of “$\mathrm{obj}[i]:$” is derived directly from the token embeddings of the language model. We propose employing simple question-answer pairs, such as “What is the obj56?” and ”Obj56 is a brown wooden shelf,” to train the projection layers $f_{\mathrm{o}}(\cdot)$ and $f_{\mathrm{a}}(\cdot)$. This approach explicitly guides the language model to establish a one-to-one correspondence between object identifiers and object-level tokens. It also aids in learning the attribute embedding and the 3D-language projection.
Through this object-level alignment, we effectively integrate the object’s individual attributes into the language model and associate them with the respective object identifiers.
3.3 Scene-level Alignment
To enable the scene understanding capability of the model, we propose the inclusion of a relation-aware token for each object. Initially, we integrate the positional information represented by $\mathbf{P}\in\mathbb{R}^{n\times 3}$ (3D coordinates) into the attribute-aware embedding through a position embedding layer $f_{\mathrm{p}}(\cdot)$. Subsequently, we employ a transformer-based relation module $r(\cdot)$ to capture the spatial relationships among objects. As a result, the relation-aware token embedding $\mathbf{F}^{\mathrm{s}}\in\mathbb{R}^{n\times d}$ can be computed by:
$\mathbf{F}^{\mathrm{s}}=r(\mathbf{F}^{\mathrm{o}}+f_{\mathrm{p}}(\mathbf{P})).$ (2)
Now we obtain the final input tokens for the $i$-th object, denoted as “$\mathrm{obj}[i]:O_{i}\ S_{i}$”, where $O_{i}$ and $S_{i}$ represent the attribute-aware and relation-aware tokens with the the token embeddings $\mathbf{F}^{\mathrm{o}}_{i}$ and $\mathbf{F}^{\mathrm{s}}_{i}$, respectively.
To learn the relation-aware token effectively, we propose using two types of data: scene-aware object captioning and scene-level question-answering, depicted as follows.
Scene-aware Object Captioning The first type is scene-aware object caption data sourced from the ScanRefer dataset. For instance, a scene-aware caption for “obj15” in Figure 2 could read: “A black tv, in the direction from the entrance and from the outside, will be on the right side of the blue curtain. On the left of the tv is a small bike.” This caption enables the model to perceive the objects in the surroundings and grasp the spatial relationships among them. To train the model, we construct single-turn QA pairs where the question asks “Provide a brief description of obj15 within the 3D scene,” and the answer is the provided caption itself.
Scene-level Question-Answering However, the original captions in the ScanRefer dataset lack object identifiers when describing the surrounding objects. This can potentially confuse our model when associating object identifiers with object tokens. Additionally, these captions mainly emphasize directional relationships, leaving the model unaware of the distances between objects. To address this, we introduce single-turn scene-level QA pairs for training. These questions encompass various types, such as “Which object is farthest from / closest to obj15?”, “List the five closest objects to obj15.”, “Which object is located at the lower / higher position, obj15 or obj34?”, and “Which object is closer to obj15, obj32 or obj56?”. The answers to these questions also include object identifiers.
For training, we combine the scene-aware caption data from ScanRefer with the proposed scene-level question-answering pairs. The projectors $f_{\mathrm{o}}(\cdot)$, $f_{\mathrm{a}}(\cdot)$, $f_{\mathrm{p}}(\cdot)$, and the relation module $r(\cdot)$ are trainable in this stage.
Table 1: Example of generated identifier-rich scene caption. The prompt for GPT-4 and the complete version of the scene caption is provided in the supplementary materials.
[Uncaptioned image] The 3D scene is a well-organized room, likely a classroom or a meeting room, with a functional layout that facilitates interaction and movement. The room features two tables (obj03 and obj04), each surrounded by a number of chairs, suggesting a collaborative workspace. The table obj03, located southeast of the whiteboard (obj33), is surrounded by chairs obj10, obj13, obj15, obj17, obj19, and obj25, while table obj04, closest to the door, is surrounded by chairs obj11, obj12, obj14, obj16, obj18, and obj20. Each table also has lamps (obj06 and obj09 for obj03, obj07 and obj08 for obj04) providing adequate lighting.
Figure 3: Example of 3D-language scene understanding tasks.
3.4 End-to-End Fine-tuning
The previous two-stage alignment process has endowed our model with robust 3D understanding and reasoning abilities. Although it can be directly employed in a zero-shot manner, we also explore the potential of fine-tuning it end-to-end, similar to previous instruction-tuning methods. In this approach, we maintain the pre-trained 3D encoder while updating the projectors, the relation module, and the language model to refine its performance on various downstream tasks.
To fully harness the potential of our model equipped with object identifiers, we create an identifier-rich scene captioning dataset, which is described in Section 4. We subsequently fine-tune our model on tasks such as 3D QA, visual grounding, and scene captioning, with additional details provided in Section 5.
4 Identifier-rich Scene Captioning Dataset
We observe that the majority of datasets within the field of 3D-language scene understanding are object-centric. In other words, these datasets primarily concentrate on an individual objects in each question-answer pair, often leading to imprecise and non-reusable language descriptions concerning the relationships between these objects and their surroundings.
To fully exploit the capabilities of our model, we generate a dataset containing comprehensive scene-level captions, and explicitly incorporating object identifiers whenever any object is referenced within the caption. We refer to this dataset as the “identifier-rich scene captioning dataset”. To create this dataset, we provide GPT-4 with ground truth information of each object, encompassing class names, 3D coordinates, and concise descriptions sourced from the ScanRefer dataset. Then we prompt it to generate identifier-rich scene captions. An example of the resulting scene caption is presented in Table 1. We generate 705 scene captions in total, each for a different scene in the ScanRefer dataset.
Through training our model on this generated dataset, we equip it with the capability to freely reference objects with identifiers when describing a complex 3D scene. Additionally, this fine-tuning process enhances the model’s overall perception and reasoning abilities within the 3D scene. In contrast to previous methods, the object identifiers in the response can be further leveraged in subsequent task flows or seamlessly visualized within the original 3D point cloud.
5 Experiments
Table 2: Evaluation results on ScanQA [3] validation set.
Method
# 3D Data
for Alignment
Backbone B-1 B-4 METEOR ROUGE CIDEr SPICE EM
VoteNet [9]+MCAN [40] - - 28.0 6.2 11.4 29.8 54.7 10.7 17.3
ScanRefer[4]+MCAN [40] - - 26.9 7.9 11.5 30.0 55.4 11.3 18.6
ScanQA [3] - - 30.2 10.1 13.1 33.3 64.9 13.4 21.0
3D-LLM (flamingo) [14] 675K VLM 30.3 7.2 12.2 32.3 59.2 - 20.4
3D-LLM (BLIP2-flant5) [14] 675K VLM 39.3 12.0 14.5 35.7 69.4 - 20.5
LLaVA (zero-shot) [21] - LLM 7.1 0.3 10.5 12.3 5.7 - 0.0
Chat-3D [33] 38K LLM 29.1 6.4 11.9 28.5 53.2 - -
Ours 38K LLM 38.4 7.3 16.1 40.1 77.1 18.8 21.1
Table 3: Evaluation results on ScanRefer [4] dataset.
Method Acc@0.25 Acc@0.5
ScanRefer [4] 37.3 24.3
MVT [15] 40.8 33.3
3DVG-Trans [45] 45.9 34.5
ViL3DRel [5] 47.9 37.7
LLM-Grounder [37] 17.1 5.3
3D-LLM (flamingo) [14] 21.2 -
3D-LLM (BLIP2-flant5) [14] 30.3 -
Ours 35.9 30.4
As shown in Figure 3, our model architecture can handle various 3D-language downstream tasks. Evaluation details on 3D QA, visual grounding, and scene captioning are provided in this section. Additional experiments can be found in the supplementary materials.
5.1 Implementation Details
For pre-training, we use ground truth annotations directly for segmenting each object, including additional attributes such as color, size, and position. To extract object features, we utilize the pre-trained Uni-3D [47] as the 3D encoder. The attribute embedding layer $f_{\mathrm{a}}(\cdot)$ and the position embedding layer $f_{\mathrm{p}}(\cdot)$ are both implemented as linear layers. The 3D-language projection module $f_{\mathrm{o}}(\cdot)$ is a three-layer MLP. The relation module $r(\cdot)$ is a one-layer transformer encoder. We choose Vicuna-7B model [7] as the LLM, which is fine-tuned from the LLaMA base model [28]. We set the base learning rate as 1e-4 for all the stage (1e-5 for relation module and LLM) with a cosine annealing schedule. The entire pre-training and fine-tuning process requires approximately 6 hours on 4 NVIDIA A40 GPUs.
Table 4: Grounding accuracy (%) on Nr3D/Sr3D [5] datasets.
Method Nr3D Sr3D
Overall Easy Hard
View
Dep
View
Indep
Overall Easy Hard
View
Dep
View
Indep
ScanRefer [4] 34.2 41.0 23.5 29.9 35.4 - - - - -
ReferIt3D [1] 35.6 43.6 27.9 32.5 37.1 40.8 44.7 31.5 39.2 40.8
3DVG-Trans [45] 40.8 48.5 34.8 34.8 43.7 51.4 54.2 44.9 44.6 51.7
MVT [15] 55.1 61.3 49.1 54.3 55.4 64.5 66.9 58.8 58.4 64.7
ViL3DRel [5] 62.6222We re-trained the model using their provided code. So the score here is different from the reported one in their paper. 70.1 55.6 59.8 64.1 72.5 74.2 68.1 58.0 73.1
Ours (fine-tuned) 34.8 42.4 27.2 33.7 35.5 39.4 43.6 29.3 37.4 39.5
Ours (refine ViL3DRel) 63.6 70.7 57.0 61.1 65.0 73.1 75.1 68.4 58.1 73.7
Table 5: Evaluation on the identifier-rich scene captioning dataset.
Method B-1 B-4 METEOR ROUGE CIDEr
Chat-3D [33] 39.5 7.5 17.7 21.1 2.2
Ours 45.6 10.1 21.9 23.4 9.7
Table 6: Ablation study on ScanQA [3] dataset. AES and OQA denote the Attribute-aware Embedding Similarity and Object-level QA components proposed in the object-level alignment phase, respectively. SOC and SQA correspond to the Scene-aware Object Caption and Scene-level QA components presented in the scene-level alignment phase, respectively.
Row Object Identifier Object Alignment Scene Alignment B-1 METEOR ROUGE CIDEr SPICE EM
AES
OQA
SOC
SQA
1
[…]
2613
[…]
2613
[…]
2613
[…]
2613
[…]
2613
23.7 8.8 20.8 36.3 8.3 8.2
2
[…]
2613
[…]
2613
[…]
2613
[…]
2613
27.8 10.7 25.5 45.9 10.6 10.9
3
[…]
2613
[…]
2613
[…]
2613
[…]
2613
26.7 11.1 26.0 46.5 11.6 10.7
4
[…]
2613
[…]
2613
[…]
2613
30.3 12.1 29.0 54.3 14.5 13.0
5
[…]
2613
[…]
2613
34.2 13.9 34.0 65.1 16.1 17.2
6
[…]
2613
33.4 14.9 36.4 69.5 17.5 16.0
7 38.4 16.1 40.1 77.1 18.8 21.1
Table 7: Ablation study of the relation-aware token.
Method B-1 MET ROU CID SPI EM
w/o relation (delete) 32.7 13.5 33.6 63.4 15.9 15.7
w/o relation (mean) 34.9 14.0 34.8 65.1 16.1 18.2
w/ relation 38.4 16.1 40.1 77.1 18.8 21.1
5.2 Evaluation on 3D Question Answering
ScanQA contains more than 40,000 question-answer pairs derived from 800 indoor scenes sourced from the ScanNet dataset. To evaluate on the ScanQA dataset, we fine-tune the pre-trained model to adapt it to the short-form answer format.
Baselines The compared baselines fall into two categories: supervised expert models and general multi-modal LLMs.
  • Supervised Expert Models: ScanQA [3] uses VoteNet [9] to generate object proposals and integrates them with language embeddings. Both ScanRefer+MCAN and VoteNet+MCAN detect 3D objects and incorporate them into a standard VQA model known as MCAN [40].
  • General Multi-modal LLMs: LLaVA [21] introduces the visual instruction tuning method, establishing a connection between the vision encoder and the LLM for general-purpose visual and language understanding. 3D-LLM [14] employs 2D VLMs like BLIP-2 [19] and Flamingo [2] as backbone, projecting 3D features into the VLM’s input space. Chat-3D [33] aims to directly align 3D scenes with LLM but is limited to object-centric tasks.
Analysis Firstly, we can categorize the caption metrics into two types: BLEU and EM, based on exact matches (sub-sequences or the entire sentence), and METEOR, ROUGE, CIDEr, and SPICE, which focus more on semantic matches. Our model exhibits lower performance on BLEU metrics compared to 3D-LLM, indicating that our model hasn’t learned the target answer format as effectively as 3D-LLM. (In the ScanQA dataset, answers are short phrases or single words.) However, our model outperforms previous methods on all other metrics, especially those based on semantic matches. This demonstrates our model’s superior scene understanding ability and its ability to respond with semantically accurate answers. Lower performance on BLEU metrics coupled with higher performance on other metrics suggests that our model’s output is more diverse and creative. This could be attributed to the language model choice and fine-tuning strategy. Further efforts should be made to achieve a better balance between diversity and exact matching.
5.3 Evaluation on 3D Visual Grounding
Our model easily addresses the 3D visual grounding task in a two-stage process. First, we detect object proposals using a pre-trained 3D segmentation model and assign them object identifiers. Second, we prompt the model to identify the object that best matches a provided description and have it respond with the corresponding object identifier. This allows us to directly obtain the segmented point cloud of the predicted object and derive its 3D bounding box, which becomes the final predicted bounding box. We evaluate our model on the ScanRefer [4] and Nr3D/Sr3D [1] datasets. For the ScanRefer dataset, we use PointGroup [16] as the 3D segmentor to be consistent with the supervised baseline ViL3DRel [5]. For the Nr3D and Sr3D datasets, we use the provided segmented annotations to keep the same with previous methods.
Baselines The compared baselines can be divided into two categories: supervised expert models and general multi-modal LLMs.
  • Supervised Expert Models: ScanRefer [4] utilizes a pretrained VoteNet backbone with a trained GRU for selecting a matching bounding box. ReferIt3D [1] employs a visio-linguistic graph neural network by enabling communication among objects. 3DVG-Trans [45] proposes a relation-aware transformer architecture to extract multi-level contextual features. MVT [15] projects the 3D scene to a multi-view space and learns position information under different views. ViL3DRel [5] designs a spatial self-attention layer that accounts for relative distances and orientations between objects.
  • General Multi-modal LLMs: LLM-Grounder [37] achieves zero-shot 3D visual grounding by utilizing an open-vocabulary visual grounding tool and using GPT-4 as the language assistant. 3D-LLM [14] predicts bounding boxes as location tokens added to the language model’s vocabularies and learned from scratch.
Analysis On the ScanRefer dataset, our model outperforms previous 3D large language models, demonstrating that our approach establishes a more robust connection between the 3D scene and LLMs. On the Nr3D/Sr3D datasets, the task involves providing the model with extracted object instances and requiring it to select one of these objects as the answer. Former 3D large language models struggle with this task because they are unable to reference a particular object from the input, while our model easily handles it with object identifiers. On both ScanRefer and Nr3D/Sr3D datasets, our model achieves comparable results with supervised baselines, indicating strong abilities in scene understanding and reasoning. Compared to supervised baselines with complex architecture designs to learn spatial relationships, our model only uses a simple relation module and achieves competitive results, demonstrating the effectiveness of our alignment method and the introduced object identifiers.
Additionally, we explore the potential of ensembling our model with powerful supervised baselines to combine their learned knowledge and achieve better performance. We observe that in a traditional two-stage grounding model, the predicted probability distribution on the object instances provides an effective filter for possible candidate answers. In most cases, these models fail when faced with several objects with high probabilities. They directly choose the one with the highest probability, while the true answer may lie among the others. In such cases, we extract up to 5 objects with a probability higher than a threshold (1e-3) and provide our model with their identifiers, allowing our model to decide on a final answer. This way, we use the knowledge of powerful baselines to filter out several most probable candidates and prompt our model to carefully consider among them. As shown in Table 4, we improve the state-of-the-art model ViL3DRel by about 1.0% in grounding accuracy through this simple model ensembling.
5.4 Evaluation on 3D Scene Captioning
We create an identifier-rich 3D scene captioning dataset with the assist of GPT-4, as explained in Section 4. We directly fine-tune our model on this dataset to show its capability of using object identifiers. In comparison, we also fine-tuned Chat-3D [33] on this dataset. Since Chat-3D doesn’t incorporate object identifiers, we excluded them from our dataset during its fine-tuning. As illustrated in  Table 5, our model, despite making some minor errors in the response, demonstrates an overall understanding of the entire scene. On the other hand, Chat-3D tends to overly focus on a specific picture attached to a wall, leading to a diminished perception of the entire scene.
5.5 Ablation Study
We conduct an ablation study on the ScanQA dataset to evaluate the effectiveness of each proposed component. The results are presented in Table 6. Firstly, we evaluate the impact of object identifiers by removing them from the language input. In Row 1, without using object identifiers and two-stage alignment, direct fine-tuning on the task resulted in extremely low performance. Row 3, compared to Row 1, indicates that only adding object identifiers to the model can enhance performance. This suggests that the language model itself can establish an initial correspondence between each object and its identifier. Row 4 to 7 systematically introduce proposed alignment strategies during pre-training. The results demonstrate that each alignment strategy contributes to the overall performance enhancement in the final model.
In addition to the ablation study of alignment methods, we delved into the effectiveness of the relation-aware token learned for each object, as shown in Table 7. Firstly, we directly delete the relation-aware token, using only an attribute-aware token for each object. Consequently, the model, unaware of any spatial relationships, performed poorly. Secondly, we attempted to replace the learned relation-aware token with the mean value of all attribute-aware tokens. In this way, each object could be aware of other objects through the relation-aware token, but without learning from the relation module, it struggled to capture complex spatial relationships, resulting in lower performance. This study highlights the effectiveness of the relation-aware token, learned by the relation module.
6 Conclusion
In this paper, we explore the potential of using object identifiers in 3D large language models. To tackle the challenges introduced by these identifiers, we propose a two-stage alignment method. This method involves incorporating the object’s attributes and spatial relationships into an attribute-aware token and a relation-aware token for each object, respectively. Our experimental results demonstrate the effectiveness of the proposed method.